# Mitigating Subgroup Disparities in Multi-Label Speech Emotion Recognition: A Pseudo-Labeling and Unsupervised Learning Approach

### v2

@@ Abstract
While subgroup disparities and performance bias are increasingly studied in computational research, fairness in categorical Speech Emotion Recognition (SER) remains underexplored. Existing methods often rely on explicit demographic labels, which are difficult to obtain due to privacy concerns. To address this limitation, we introduce an Implicit Demography Inference (IDI) module that leverages pseudo-labeling from a pre-trained model and unsupervised learning using k-means clustering to mitigate bias in SER. Our experiments show that pseudo-labeling IDI reduces subgroup disparities, improving fairness metrics by over 33% with less than a 3% decrease in SER accuracy. Also, the unsupervised IDI yields more than a 26% improvement in fairness metrics with a drop of less than 4% in SER performance. Further analyses reveal that the unsupervised IDI consistently mitigates race and age disparities, demonstrating its potential in scenarios where explicit demographic information is unavailable.
@@ 1 Introduction
Speech Emotion Recognition (SER) is essential in human–computer interaction, enhancing user experiences across various applications [1]. As SER becomes increasingly integral to applications ranging from virtual assistants to therapeutic tools, ensuring that these systems operate fairly and without bias is more critical than ever [2, 3]. However, conventional debiasing methods predominantly depend on explicit demographic annotations (e.g., gender, age, race), which are not only challenging to collect due to privacy concerns and ethical issues but are also notably scarce in real-world data [4, 5, 6, 7, 8].
@@
Our experiments on the CREMA-D database [13] provide compelling evidence for the efficacy of this approach. We observe that pseudo-labeling yields improvements in key fairness metrics by over 33%, while the unsupervised clustering approach enhances these metrics by approximately 26%, all with only a minor compromise in overall SER accuracy. Furthermore, the SER model debiased using the unsupervised clustering method can also suppress age and race bias, proving the generalizability of our method. These results underscore the potential of the IDI module to serve as a robust and scalable solution for debiasing in SER, especially in environments where explicit demographic data is limited.
Figure 1: Illustration of proposed debias framework.
@@ 3 Methodology
In this work, we propose a novel framework to mitigate subgroup disparities in gender in multi-label SER without relying on explicit demographic annotations. Central to our approach is the Implicit Demography Inference (IDI) Module, which infers group labels from speech data using a combination of pseudo-labeling and unsupervised clustering. These inferred group labels are then utilized to guide a debias training process. An overview of the proposed framework is illustrated in Fig. 1. In the following sections, we detail the architecture of our proposed IDI module and its integration with debias training techniques.
@@ 3.1 Implicit Demography Inference (IDI) Module
Unsupervised Clustering: We extract embeddings using the ECAPA-TDNN model [12] and apply K-Means clustering [19] (details about parameters are in Appendix Table A1) to these embeddings. The cluster assignments derived from this process are treated as group labels, reflecting latent structures within the data that may correspond to demographic differences. We set the cluster sizes to 2, 4, 8, 16, and 32.
@@ 3.2 Debias Training with Inferred Group Labels
Once the group labels are inferred by the IDI Module, they are incorporated into the debias training of the SER model. Specifically, the inferred labels partition the training data into subgroups, allowing us to impose fairness constraints that decrease performance disparities across these groups. We explore four debiasing methods that leverage these inferred group labels:
Reweighting (RW) [20]: This approach adjusts the importance of each sample based on the frequency of its subgroup within each emotion class. The overall weighted loss is computed as:
@@
where $\lambda_{GD}$ is a hyperparameter and $n_{g}$ is the group size. We set $\lambda_{GD}$ to 4.
@@ 4.2 Dataset
Table 1: Summary of results in SER accuracy (F1 and ACC) and fairness (TPRgap and DPgap) on gender. Super. specifies the level of supervision for debias. indicates debias with ground truth gender labels. represents the pseudo-labeling approach, and refers to bias unsupervised methods. For unsupervised IDI, we use k-means with k=16. Overall averages the four IDI methods. The Gain represents the improvement compared to the ERM method. $\uparrow$ indicates that higher values correspond to better performance.
@@ 5.1 Performance and Effects of Pseudo-labeling
Our results show that pseudo-labeling reduces subgroup disparities in multi-label SER. As seen in Table 1, using pseudo-labels from a pre-trained gender detection model improves the TPRgap fairness metric by 33% and DP by 42% relative to the ERM baseline, with less than a 3% drop in overall accuracy. Compared to PGDRO [27], our modified reweighting strategy with pseudo-labels achieves a better balance between fairness and performance.
@@ 5.2 Performance and Effects of Unsupervised Learning
The number of clusters plays a critical role in the effectiveness of this method. Using too few clusters may result in overly coarse groupings that fail to capture the underlying nuances in the data, while employing too many clusters can fragment the data and introduce noise, leading to less meaningful subgroup distinctions. We found that a cluster size of 16 provides a balanced granularity, offering significant improvements in fairness metrics (26% in TPRgap and 27% in DPgap) while maintaining a minor drop in overall performance (a 4% decrease in F1 and a 3% decrease in ACC), as shown in Table 1.
@@
Figure 2: Subfigure at left presents the TPRgap while subfigure at right shows the Hamming Accuracy. Bars show the performance of the proposed unsupervised debiasing method evaluated across four techniques under various numbers of clusters.
@@
Figure 4: Illustration of samples in Fold 5 using t-SNE (details about parameters are in Appendix Table A1), with ECAPA-TDNN embeddings and a cluster size of 16. Circles and triangles represent male and female speakers, respectively, while different colors indicate different clusters.
@@ 5.3 Potential of Unsupervised Clustering
Table 2 presents overall performance across subgroups, demonstrating that the unsupervised method with K=16 clusters improves TPRgap metric by over 12.75%, while maintaining an SER performance decrease of less than 4.76% and 3.55% in F1 and ACC, respectively. This suggests that the method’s performance is robust across different demographic attributes, highlighting its potential for broad applicability for subgroup disparities in multi-label SER tasks.
@@ 7 Conclusion and Future Work
This work empirically examines various debiasing methods in supervised, pseudo-label, and unsupervised learning to mitigate subgroup disparities without explicit demographic information. We leverage pseudo-labels generated by a pre-trained model and introduce an unsupervised clustering approach that integrates with existing debiasing techniques originally designed for demographic labels. Our findings demonstrate that pseudo-label learning improves TPR and DP metrics by 33.33% and 47.90%, respectively, with only a minor 3.02% degradation in SER accuracy. Similarly, our proposed unsupervised method achieves 26.44% and 26.54% improvements in TPR and DP metrics, respectively, with just a 4.22% drop in SER performance. For future work, we will explore more possible bias attributes like content [32], language [33], or speaking style [34].
@@
The Supplementary Material provides further details, including Table A1, which outlines the functions for K Means clustering and t-SNE plotting. Additionally, Table A2 presents the data distribution at a 1:20 ratio for Fold 1 of the CREMA-D emotion database, and Table A3 details the distribution of emotions, gender, race, and age across all five folds of the database.
@@ Appendix A Implementation Details
We use the Adam optimizer with a 0.0001 learning rate. The batch size and epoch are set as 32 and 50, respectively. We choose the best models according to the lowest value of the class-balanced cross-entropy loss on the development set. We use two Nvidia Tesla v100 GPUs with 32 GB memory. The total number of GPU hours is around 500.
Table A1: Scikit-learn version is 1.2.2
K Means t-SNE
Parameter Value Parameter Value
init greedy k-means++ init pca
n_clusters [2,4,8,16,32] method barnes_hut
random_state 42 random_state 42
max_iter 1000 n_components 2
batch_size 32 perplexity 40
compute_labels TRUE metric euclidean
reassignment_ratio 0.01 n_iter deprecated
tol 1.00E-04 n_iter_without_progress 300
algorithm lloyd learning_rate auto

### v3

@@ Abstract
While subgroup disparities and performance bias are increasingly studied in computational research, fairness in categorical Speech Emotion Recognition (SER) remains underexplored. Existing methods often rely on explicit demographic labels, which are difficult to obtain due to privacy concerns. To address this limitation, we introduce an Implicit Demography Inference (IDI) module that leverages pseudo-labeling from a pre-trained model and unsupervised learning using k-means clustering to mitigate bias in SER. Our experiments show that pseudo-labeling IDI reduces subgroup disparities, improving fairness metrics by over 28% with less than a 2% decrease in SER accuracy. Also, the unsupervised IDI yields more than a 4.6% improvement in fairness metrics with a drop of less than 3.6% in SER performance. Further analyses reveal that the unsupervised IDI consistently mitigates race and age disparities, demonstrating its potential when explicit demographic information is unavailable.
@@ 1 Introduction
Speech Emotion Recognition (SER) is essential in human–computer interaction, enhancing user experiences across various applications [1]. As SER becomes increasingly integral to applications ranging from virtual assistants to therapeutic tools, ensuring that these systems operate fairly and without bias is more critical than ever [2, 3]. However, conventional debiasing methods predominantly depend on explicit demographic annotations (e.g., gender, age, race), which are not only challenging to collect due to privacy concerns and ethical issues but are also notably scarce in real-world data [4, 5, 6, 7, 8, 9, 10].
@@
Our experiments on the CREMA-D database [15] provide compelling evidence for the efficacy of this approach. We observe that pseudo-labeling yields improvements in key fairness metrics by over 28%, while the unsupervised clustering approach enhances these metrics by approximately 4.6%, all with only a minor compromise in overall SER accuracy. Furthermore, the SER model debiased using the unsupervised clustering method can also suppress age and race bias, proving the generalizability of our method. These results underscore the potential of the IDI module to serve as a robust and scalable solution for debiasing in SER, especially in environments where explicit demographic data is limited.
Figure 1: Illustration of the proposed debias framework.
@@ 3 Methodology
We propose a novel framework to mitigate subgroup disparities in gender in multi-label SER without relying on explicit demographic annotations. Central to our approach is the Implicit Demography Inference (IDI) Module, which infers group labels from speech data by combining pseudo-labeling and unsupervised clustering. These inferred group labels are then utilized to guide a debias training process. An overview of the proposed framework is illustrated in Fig. 1. In the following sections, we detail the architecture of our proposed IDI module and its integration with debias training techniques.
@@ 3.1 Implicit Demography Inference (IDI) Module
Unsupervised Clustering: We extract embeddings using the ECAPA-TDNN model [14] and apply K-Means clustering [21] (details about parameters are in Appendix Table LABEL:tab:appendix_sklearn) to these embeddings. The cluster assignments derived from this process are treated as group labels, reflecting latent structures within the data that may correspond to demographic differences. We set the cluster sizes to 2, 4, 8, 16, and 32.
@@ 3.2 Debias Training with Inferred Group Labels
Once the IDI Module infers the group labels, they are incorporated into the debias training of the SER model. Specifically, the inferred labels partition the training data into subgroups, allowing us to impose fairness constraints that decrease performance disparities across these groups. We explore four debiasing methods that leverage these inferred group labels:
Reweighting (RW) [22]: This approach adjusts the importance of each sample based on the frequency of its subgroup within each emotion class. The overall weighted loss is computed as follows:
@@
where $\lambda_{GD}$ is a hyperparameter and $n_{g}$ denotes the group size. In our experiment, we set $\lambda_{GD}$ to 4.
@@ 4.2 Dataset
Table 1: Summary of results in SER accuracy (F1 and ACC) and fairness (TPRgap and DPgap) on gender. Super. specifies the level of supervision for debias. indicates debias with ground truth gender labels. represents the pseudo-labeling approach, and refers to bias unsupervised methods. For unsupervised IDI, we use k-means with k=16. Overall averages the four IDI methods. The Gain represents the improvement compared to the ERM method. $\uparrow$ indicates that higher values correspond to better performance; $\downarrow$ indicates the opposite.
@@ 5.1 Performance and Effects of Pseudo-labeling
Our results show that pseudo-labeling reduces subgroup disparities in multi-label SER. As seen in Table 1, using pseudo-labels from a pre-trained gender detection model improves the TPRgap fairness metric by 28.78% and DP by 29.13% relative to the ERM baseline, with less than a 1.5% drop in overall accuracy. Compared to PGDRO [29], our modified reweighting strategy with pseudo-labels achieves a better balance between fairness and performance.
@@ 5.2 Performance and Effects of Unsupervised Learning
The number of clusters plays a critical role in the effectiveness of this method. Using too few clusters may result in overly coarse groupings that fail to capture the underlying nuances in the data, while employing too many clusters can fragment the data and introduce noise, leading to less meaningful subgroup distinctions. We found that a cluster size of 16 provides a balanced granularity, offering significant improvements in fairness metrics (20.05% in TPRgap and 4.61% in DPgap) while maintaining a minor drop in overall performance (a 4.76% decrease in F1 and a 3.55% decrease in ACC), as shown in Table 1.
@@
Figure 2: The left subfigure presents the TPRgap, while the right one depicts the Hamming accuracy. The bars illustrate the performance of the proposed unsupervised debiasing method, evaluated across four different techniques with varying cluster sizes.
@@
Figure 4: Illustration of samples in Fold 5 using t-SNE (details about parameters are in Appendix Table LABEL:tab:appendix_sklearn), with ECAPA-TDNN embeddings and a cluster size of 16. Circles and triangles represent male and female speakers, respectively, while different colors indicate different clusters.
@@ 5.3 Potential of Unsupervised Clustering
Table 2 presents overall performance across subgroups, demonstrating that the unsupervised method with K=16 clusters improves TPRgap metric by over 12.75%, while maintaining an SER performance decrease of 4.76% and 3.55% in F1 and ACC, respectively. This suggests that the method’s performance is robust across different demographic attributes, highlighting its potential for broad applicability for subgroup disparities in multi-label SER tasks.
@@ 7 Conclusion and Future Work
This work empirically examines various debiasing methods in supervised, pseudo-label, and unsupervised learning to mitigate subgroup disparities without explicit demographic information. We leverage pseudo-labels generated by a pre-trained model and introduce an unsupervised clustering approach that integrates with existing debiasing techniques originally designed for demographic labels. Our findings demonstrate that pseudo-label learning improves TPR and DP metrics by 28.78% and 29.13%, respectively, with only a minor 4.38% degradation in SER accuracy. Similarly, our proposed unsupervised method achieves 20.05% and 4.61% improvements in TPR and DP metrics, respectively, with just a 4.76% drop in SER performance. For future work, we will explore more possible bias attributes like content [34], language [35], or speaking style [36].
8 Acknowledgement
We thank Taiwan’s National Center for High-Performance Computing (NCHC) at the National Applied Research Laboratories (NARLabs) for providing the computational and storage resources.
@@ 8 Acknowledgement
The Supplementary Material furnishes a comprehensive account of our experimental setup. Table A1 and A2 specify the implementations used for K-Means clustering and t-SNE visualization, respectively. Additionally, Table A3 presents the data distribution at a 1:20 ratio for Fold 1 of the CREMA-D emotion database, and Table A4 details the distribution of emotions, gender, race, and age across all five folds of the database.
@@ Appendix A Implementation Details
We use the Adam optimizer with a 0.0001 learning rate. The batch size and epoch are set as 32 and 50, respectively. We choose the best models according to the lowest value of the class-balanced cross-entropy loss on the development set. We use two Nvidia Tesla V100 GPUs with 32 GB of memory. The total number of GPU hours is around 500.
Table A1: Detailed configuration of the K-Means clustering algorithm as implemented in scikit-learn v1.2.2. This table lists every parameter we controlled.
Parameter Value
Initialization k-means++ (greedy)
Number of clusters [2, 4, 8, 16, 32]
Random seed 42
Maximum iterations 1 000
Mini-batch size 32
Compute final labels True
Reassignment ratio 0.01
Convergence tolerance $1\times 10^{-4}$
Algorithm Lloyd’s method
Table A2: Detailed configuration of the t-SNE embedding parameters used with scikit-learn v1.2.2.
Parameter Value
Initialization PCA
Neighbor-search method Barnes–Hut
Random seed 42
Target dimensionality 2
Perplexity 40
Distance metric Euclidean
Iterations without progress before abort 300
Learning rate auto